OpenStudy (anonymous):

Will medal and fan all who help Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B. How fast did Car A travel?

OpenStudy (anonymous):

55mph

OpenStudy (anonymous):

let x=speed of car B x+22=speed of car A distance=travel time*speed(same for both cars) 4x=2.4(x+22)

OpenStudy (anonymous):

@H3LPN33DED wuts ur final answer ?

OpenStudy (anonymous):

You have the formula, Velocity=distance/time hence we'll use that to solve for the velocity of both Cars A and B since Car A traveled for 2.4 hours, then we'll have Va = d/2.4 -----equation 1 since Car B traveled for 4 hours, then, Vb = d/4 -----equation 2 NOTE that the distance traveled by Cars A and B are equal getting the distance from equations 1 & 2 Va = d/2.4 d=2.4Va -------equation 3 Vb = d/4 d=4Vb -------equation 4 since distance of A and B are equal then we can equate equations 3&4 2.4Va=4Vb ------equation 5 The problem states that Car A traveled 22mph faster than car B. Therefore we'll have the equation as: Va=Vb+22 -------equation 6 substituting equation 6 to equation 5 we'll have 2.4Va=4Vb 2.4(Vb+22)=4Vb 2.4Vb +22(2.4)=4Vb 4Vb-2.4Vb=52.8 1.6Vb=52.8 Vb=33mph substituting Vb to equation 6 Va=Vb+22 Va=33+22 Va=55mph

OpenStudy (anonymous):

r-22=33 r=55

OpenStudy (anonymous):

@ah123456x ohhhhhh I gotcha ty